Boardwalk Properties owns and operates a strip mall with several vacancies. In an effort to revitalize the strip mall, they enter into negotiations with four major businesses to move into the mall. Each of them states that they are interested, but they are not interested unless Boardwalk is able to bring in a major anchor tenant. (An anchor tenant is a large tenant at a shopping center whose presence attracts other tenants and shoppers.) To this end, Boardwalk approaches Bullseye, a major big-box retailer, to be the anchor tenant. After negotiations, they enter into a contract, wherein Bullseye acknowledges that it is an anchor tenant and that its presence will affect other tenants in the area. To that end, it is to pay $7500 a month in rent for at least five years. This is a twenty-percent discount from the fair rental value for the space. After Bullseye signs the agreement with Boardwalk, the four previously mentioned businesses sign leases with Boardwalk.

Two years into the lease, Bullseye approaches Boardwalk and attempts to negotiate decreased rent. It tells Boardwalk that, unless it reduces rent to $6500, it will leave the strip mall and move into the new shopping center on the other side of town. Boardwalk refuses, and Bullseye abandons the premises three months later. Over the course of the following six months, the four other tenants previously mentioned leave. Boardwalk is able to find a tenant to take the space previously occupied by Boardwalk, but only by offering to accept $6000 a month in rent.

Boardwalk sues Bullseye, seeking (1) lost rent from Bullseye's broken lease, (2) lost rent from the broken leases of the other four tenants, and (3) punitive damages for intentional breach of contract. To what is Boardwalk entitled from Bullseye? It is not necessary that you calculate the exact amount of money owed to Boardwalk, but (1) you should address each element of damages Boardwalk is seeking, (2) to the extent that you can, you should explain how to calculate the lost damages.

Please list the element of contract law that applied to each situation.
Answer #1

In some states in USA, like New York, the lessor can even choose not to find any new lessee and claim the damages completely from the previous occupant. So, Broadwalk can sue Bulleye for Breach of Lease. The method to calculate the damages is:

Damages for three years:

12 (months)* 7500 (Lease amount) *3 ( No. of remaining years)= 12 * 7500 * 3= $ 270,000

In case where Broadwalk has found a now occupant for the property, they can claim the opportunity cost i.e.

$ 7500- $ 6000 = $ 1500 per month.

The remaining calculation would be the same i.e. 12 * 1500 * 3 = $ 54,000